Be sure to answer all parts. A 0.9340 g sample of a mixture of NaCl and KCl is dissolved in water, and the solution is then treated with an excess of AgNO3 to yield 1.913 g of AgCl. Calculate the percent by mass of each compound in the mixture.

% mass NaCl

% mass KCl
Answer #1

Lets take mass of Nach m(Nad) = x H & mass of kcl mckelly mp_ Wall & Ag No → Agd + NaNO₃ Number of - Number of moles of Nae mData 2:2283 - 2:45y + 1.924 = 1.913 2.2.883 -1.913 = 2.454 -1.92y 0:3753 = 0.53y y = 0.3753 /0.53 y 20.7081 x= 0.9 340-y x =